OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x – 5| + 2 = 5 Step 2: |x – 5| = 5 – 2 Step 3: |x – 5| = 3 Which of the following is a correct next step to solve the equation? x + 5 = -3 -x – 5 = 3 x - 5 = -3 x + 5 = 3

OpenStudy (anonymous):

i sowwy idk

OpenStudy (anonymous):

its alrighttt

OpenStudy (jdoe0001):

\(\bf |x - 5| + 2 = 5\implies |x - 5|=5-2\implies |x - 5|=3\\ \quad \\ |x - 5|=3\implies \begin{cases} +(x - 5)=3\\ \quad \\ \bf -(x - 5)=3 \end{cases}\)

OpenStudy (jdoe0001):

absolute value expressions will give you 2 values, one for each scenario, for "x" so solve each linear equation for "x" to get both values

OpenStudy (anonymous):

i would solve each linear equation for x to get the values if i knew what it was can you please give me the answer

OpenStudy (anonymous):

@jdoe0001 ima stoner, ima stonerr, ima stonerrr

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=zNtqneXTSWM

OpenStudy (anonymous):

aint nobdy got time for that